# Towards Language-Augmented Multi-Agent Deep Reinforcement Learning

### v1

@@ Abstract
Communication is a fundamental aspect of coordinated behavior in multi-agent reinforcement learning. Yet, most prior works in this field have focused on emergent communication protocols developed from scratch, often resulting in inefficient or non-interpretable systems. Inspired by the role of language in natural intelligence, we investigate how grounding agents in a human-defined language can improve learning and coordination of multiple embodied agents. We propose a framework in which agents are trained not only to act but also to produce and interpret natural language descriptions of their observations. This language-augmented learning serves a dual role: enabling explicit communication between agents and guiding representation learning. We demonstrate that agents trained with our method outperform traditional emergent communication baselines across various tasks. Our analysis reveals that language grounding leads to more informative internal representations, better generalization to new partners, and improved capability for human-agent interaction. These findings demonstrate the effectiveness of integrating structured language into multi-agent learning and open avenues for more interpretable and capable multi-agent systems.
@@ 1 Introduction
Multi-agent deep reinforcement learning (MADRL) studies how coordinated behavior can be learned by groups of agents to optimize task performance. One crucial skill for facilitating social interactions is communication, as observed in many social animal species [45, 43], and, of course, humans [11]. It enables a great range of social abilities: from sharing information about local observations to negotiation, knowledge transmission, and teaching. Inspired by these natural systems, artificial agents can benefit from learning to communicate, improving coordination and overall group performance.
@@
Following these observations, a promising avenue for artificial intelligence is in the investigation of how language can help artificial agents when learning to behave in complex environments and social settings. Many research works have pursued this direction in the context of single-agent reinforcement learning, using language to guide training by describing the world [18, 3], providing and exploring high-level goals [6], and decomposing tasks [19, 53]. However, in the multi-agent domain, the learning of language for guiding embodied agents and communicating in social settings is still an understudied subject. On one side, emergent communication approaches have dominated MADRL research, having agents develop their own communication system from scratch in the course of training [56]. While this can be effective, it often leads to inefficient communication systems [4, 10] and lacks the guidance provided by a pre-defined set of symbols and rules to describe the world and share information with anyone. Some works have studied learning to communicate with natural language [26, 13], but very few of them do so in embodied settings where agents explore their environment through physical interactions. Finally, language is also at the center of the recent trend focused on using large language models (LLMs) in various social settings [1, 36]. While this direction offers many promising results, it focuses on using pre-trained language models, thus missing the insights on how language learning and behavior learning interact.
@@
Through extensive experiments, we study the many advantages provided by this language-augmented training scheme. Language-augmented agents are shown to outperform emergent communication baselines across multiple tasks. We provide various results indicating that language training helps agents build better representation, including embedding visualizations and ablation studies. Further experiments showcase multiple benefits of language-augmented multi-agent systems for generalizing to new partners and interacting with humans.
@@ 4.2 Agent Architecture
To design language-augmented agents, we build a neural network architecture, shown in Figure 1, that integrates the tools required for learning to (i) complete a given embodied task and (ii) use and understand a given language. Problem (i) is addressed using classical elements of MADRL. Each agent $i$ is equipped with a local behavior policy $\pi^{i}(o$, and a centralized value function $V^{i}$, as described in Section 3. Prior to computing the policy and value, the local and joint observations are passed through encoders – "Obs Encoder" and "Joint-Obs Encoder", respectively – to extract valuable information. The two observation encoders are defined as recurrent neural networks to allow some memorization of information gathered in previous time steps. With only these four modules, an agent can learn to behave in a multi-agent setting, but has no communication capacities.
@@ Appendix D Details on experimental settings
[…]
@@
[…]

### v2

@@ Abstract
Most prior works on communication in multi-agent reinforcement learning have focused on emergent communication, which often results in inefficient and non-interpretable systems. Inspired by the role of language in natural intelligence, we investigate how grounding agents in a human-defined language can improve the learning and coordination of embodied agents. We propose a framework in which agents are trained not only to act but also to produce and interpret natural language descriptions of their observations. This language-augmented learning serves a dual role: enabling efficient and interpretable communication between agents, and guiding representation learning. We demonstrate that language-augmented agents outperform emergent communication baselines across various tasks. Our analysis reveals that language grounding leads to more informative internal representations, better generalization to new partners, and improved capability for human-agent interaction. These findings demonstrate the effectiveness of integrating structured language into multi-agent learning and open avenues for more interpretable and capable multi-agent systems. [^1] Our code is available at github.com/MToquebiau/LAMARL
@@ 1 Introduction
Multi-agent deep reinforcement learning (MADRL) studies how groups of artificial agents can be trained to optimize their performance on a given task. One crucial skill for facilitating social interactions is communication, as observed in many social animal species [46, 44], and, of course, humans [11]. It enables a great range of social abilities: from sharing information about local observations to negotiation, knowledge transmission, and teaching. Inspired by these natural systems, artificial agents can benefit from learning to communicate, improving coordination and overall group performance.
@@
Following these observations, a promising avenue for artificial intelligence is in the investigation of how language can help artificial agents when learning to behave in complex environments and social settings. Many research works have pursued this direction in the context of single-agent reinforcement learning, using language to guide training by describing the world [18, 3], providing and exploring high-level goals [6], and decomposing tasks [19]. However, language-augmented training in multi-agent settings is still an understudied subject. The dominating approach in MADRL research is emergent communication, where agents develop their own communication system from scratch in the course of training [56]. While this can be effective, it often leads to inefficient communication systems [4, 10] and lacks the guidance provided by a pre-defined set of symbols and rules to describe the world and share information with anyone. Some works have studied learning to communicate with natural language [26, 13], but very few of them do so in embodied settings where agents explore their environment through physical interactions. Finally, language is also at the center of the recent trend focused on using large language models (LLMs) in various social settings [1, 37]. While this direction offers many promising results, it focuses on using pre-trained language models, thus missing the insights on how language learning and behavior learning interact.
@@
Through extensive experiments, we study the many advantages provided by this language-augmented training scheme. Language-augmented agents are shown to outperform emergent communication baselines across multiple tasks. We provide various results indicating that language training helps agents build better representation, including embedding visualizations and ablation studies. Further experiments showcase multiple benefits of language-augmented multi-agent systems for generalizing to new partners and interacting with humans. Finally, we provide a discussion on the costs, benefits and various potentials of this approach, and argue for more works in this direction.
@@ 4.2 Agent Architecture
To design language-augmented agents, we build a neural network architecture, shown in Figure 1, that integrates the tools required for learning to (i) complete a given embodied task and (ii) use and understand a given language. Problem (i) is addressed using classical elements of MADRL. Each agent $i$ is equipped with a local behavior policy $\pi^{i}(o^{i}_{t})$, and a centralized value function $V^{i}(\mathbf{o_{t}})$, as described in Section 3. Prior to computing the policy and value, the local and joint observations are passed through encoders – "Obs Encoder" and "Joint-Obs Encoder", respectively – to extract valuable information. The two observation encoders are defined as recurrent neural networks to allow some memorization of information gathered in previous time steps. With only these four modules, an agent can learn to behave in a multi-agent setting, but has no communication capacities.
@@ Appendix D Details on experimental settings
$V=\left\{\text{Prey, North, South, East, West, Center}\right\}.$
@@
$V=\left\{\text{North, South, East, West, Center, Red, Green, Blue}\right\}.$